if a = log 3 x 2 + (log 3 x) 2 – 10 and x a = 1/x 2 then x = ? a.

if a = log3x+ (log3 x) – 10 and xa  = 1/x then x = ?
a.

Grade:9

1 Answers

Shanmukha
8 years ago
From  xa  = 1/x2
         xa+2=1
        a+2=0 
we get a = -2
Now put log3 x=t and sub. a=-2 in the given equation.
You will get a quadratic equation. Find roots and sub. t and find x
ANs: x = 1/81 or 9